A person can throw a ball at a maximum horizontal range of 120m. How high would the ball go straight up? [6 marks]

1
Expert's answer
2022-06-02T10:25:46-0400

"\theta=45^\circ"

"R=120\:\rm m"

"R=\frac{v_0^2\sin2\theta}{g}=\frac{v_0^2}{g}"

"h_{\max}=\frac{v_0^2\sin^2\theta}{2g}=\frac{R}{2}\sin^2\theta"

"h_{\max}=\frac{120\;\rm m}{2}\sin^245^\circ=30\:\rm m"